OpenStudy (anonymous):

2k-3c=6 6k-9c=9 (solving systems by using elimination)

OpenStudy (anonymous):

You can think of the method of elimination much like subtraction. We can multiply the first equation by 3 in order to get 6k-9c=18 and then subtract the second from the first in order to get: 6k-9c=18 -6k-9c=9 Which results in 0=9. We have thus arrives at, I think, the point of the problem. Since this statement can't possibly be true I believe the solution is that these two lines do not intersect. Rather they are parallel.